OpenStudy (kmesbah):

How is the slope of the tangent line is lim delta f over delta x as delta x approaches 0 and delta f is fixed? I guess the line would tend to be vertical as delta x approaches to 0 in this case. So could you please clear that out for me? Thank you!

OpenStudy (phi):

the slope is \[ \lim_{\Delta x \rightarrow 0} \frac{f(x+\Delta x ) - f(x)}{\Delta x }\] using h (easier to type), think of two points at (x, f(x) ) and (x+h, f(x+h) ) where h is tiny (very tiny) slope is change in y divided by change in x (here f(x) is the y value) \[ m = \frac{f(x+h) - f(x)}{(x+h) - x} = \frac{f(x+h) - f(x)}{h}\] if we let h approach (but never reach 0), we write: \[f'(x)= \lim_{h \rightarrow 0} \frac{f(x+h) - f(x)}{h}\] which is "like the slope", but for two points that are infinitesimally close to each other. By definition, this is the derivative of f(x)

OpenStudy (kmesbah):

So basically delta f is not fixed because obviously it is equal to \[f (x _{0} + \Delta x) - f(x_{0})\]. And as delta x approaches 0 delta f also approaches 0. Thank you!